solve (D³ – 7DD'² –6D'³)z = sin(x+2y) + e^2x+y
1
Expert's answer
2020-11-09T19:52:19-0500
"Solution"

"m^3-7m =0;\ m(m^2-7)=0\\
\therefore m_1=0,m_2=\sqrt{7}\\
C.F=\phi_1 (y+m_1 \cdot x)+\phi_2 (y+m_2 \cdot x)+x\ \phi_3 (y+m_2 \cdot x)\\
C.F=\phi_1 (y+0 \cdot x)+\phi_2 (y+(\sqrt{7}) \cdot x)+x\ \phi_3 (y+(\sqrt{7}) \cdot x)\\"

"P.I=\frac{sin(x+2y) + e^{2x}+y}{D³ – 7DD'² –6D'³}\\
F(D,D')=D³ – 7DD'² –6D'³;\ Where\ a=1, b=2\\
f(a,b)=(1,2)=1^3-7(1)(2)^2-6(2)^3=1-7(4)-6(8)\\
\implies-75 \ne0\\"

"P.I=\frac{1}{f(a,b)}\intop \intop sin\ v\ \delta v \delta v + \intop \intop (e^{2x}+y)\ \delta y \delta x \\
\implies \frac11\intop \intop -cos\ v\ \delta v \delta v \\
\implies 1\intop \intop -sin\ v\ \delta v =cot\ v= cot(x+2y)\\
and\\
\implies \intop \intop (e^{2x}+y)\ \delta y \delta x =\intop e^{2x}+y)\ \delta y = e^{2x}y+\frac{y^2}{2}\\
\implies \frac{y^2}{2}x+y \frac12 e^{2x}\\
P.I =cot(x+2y)+\frac{y^2}{2}x+y \frac12 e^{2x}\\"

General Solution

"C.F+P.I= \phi_1 (y+0 \cdot x)+\phi_2 (y+(\sqrt{7}) \cdot x)+x\ \phi_3 (y+(\sqrt{7}) \cdot x)+cot(x+2y)+\frac{y^2}{2}x+y \frac12 e^{2x}\\"